What is the total land area in Europe as a percent of Earth?
Europe has 3800 and Earth has 57,900. do i divide or what? i am lost
thank you

Answer by stanbon(48541)

What is the total land area in Europe as a percent of Earth?
Europe has 3800 and Earth has 57,900.
-------------
Percent means you have a denominator of 100.
--------------
Your Problem:
3800/57900 = x/100
Cross-multiply to get:
x = 380000/57900
x = 6.56
---------
The answer is 6.56%
=====================
Cheers,
Stan H.